# Astrophysical Consequences of Black Holes in Asymptotically Safe Quantum Gravity: Shadows and Strong Gravitational Lensing

### v2

@@ IV Strong gravitational lensing and its observables
In this section, we examine the phenomenon of strong gravitational lensing in the vicinity of a black hole in asymptotically safe quantum gravity. Our goal is to understand how the black hole parameters $S_{0}$, $S_{2}$, $m_{0}$, and $m_{2}$ influence the observable outcomes of strong lensing and to compare the lensing behavior of the asymptotically safe quantum gravity black hole with that of the Schwarzschild black hole ($S_{0}=0=S_{2}$).
We begin our analysis by considering the strong deflection angle of a light ray within the equatorial plane ($\theta=\frac{\pi}{2}$) of the black hole.
IV.1 The strong lensing angle
To obtain the strong deflection angle of a light ray in the equatorial plane ($\theta=\frac{\pi}{2}$) of a black hole, we rewrite first the metric (2) by redefining the quantities $r$, $t$, $S_{0}$, $S_{2}$, $m_{0}$ and $m_{2}$ in units of the radius $2M$, so that $t\rightarrow t/2M$, $r\rightarrow r/2M$ ,$S_{0}\rightarrow S_{0}/2M$,$S_{2}\rightarrow S_{2}/2M$,$m_{0}\rightarrow 2Mm_{0}$ and $m_{2}\rightarrow 2Mm_{2}$, respectively. Hence, the asymptotically safe quantum geometric metric (2) takes in the equatorial approximation the form
@@ IV.1 The strong lensing angle
The unstable circular photon orbit, with radius $r_{ph}$, can be obtained from the conditions on the effective potential, given by $\left.\frac{dV_{eff}}{dr}\right|_{r_{ph}}=0$ and $\left.\frac{d^{2}V_{eff}}{dr^{2}}\right|_{r_{ph}}<0$, respectively. Thus, the radius of the photon sphere $r_{ph}$ is the largest real root of the equation [115, 116, 117]:
@@
The behavior of the photon sphere radius, $r_{ph}$, and the critical impact parameter, $u_{ph}$, are expressed as functions of the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$ (see Figs. 3 and 4). It is observed from Fig. 3 that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase while keeping the other parameters fixed, the photon sphere radius $r_{ph}$ decreases. Conversely, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters held constant, increases the photon sphere radius, $r_{ph}$. Furthermore, the photon sphere radius $r_{ph}$ reaches its maximum ($r_{ph}=1.5$ ) when both $S_{0}\rightarrow 0$ and $S_{2}\rightarrow 0$,corresponding to Schwarzschild black hole.
It is also observed from Fig. 4 that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase while keeping the other parameters fixed, the critical impact parameter, $u_{ph}$, decreases. Conversely, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters held constant, results in an increase in the critical impact parameter, $u_{ph}$ (see Table 2). Furthermore, the critical impact parameter $u_{ph}$ reaches its maximum ($u_{ph}=2.59808$ ) when both $S_{0}\rightarrow 0$ and $S_{2}\rightarrow 0$, corresponding to Schwarzschild black hole.
@@
The strong deflection angle in the asymptotically safe quantum gravity black hole spacetime can be obtained as a function of the closest approach distance $r_{0}$ as [31]
@@
The strong deflection angle $\alpha_{D}(r_{0})$ depends on the relationship between $r_{0}$ and $r_{ph}$, and it increases when these two values are approximately equal (i.e., $r_{0}\approx r_{ph}$). Hence, we introduce a new variable z defined as [118, 119]
@@
where the integral must be evaluated numerically. We numerically calculate the values of the critical impact parameter $u_{ph}$ and of the strong lensing coefficients $\bar{a}$ and $\bar{b}$ at $r=r_{ph}$ for the asymptotically safe quantum gravity black hole parameters $S_{0}=0,0.5$,$S_{2}=0,0.1,0.2,0.3$, $m_{0}=0.5,1$, and $m_{2}=0.3,0.6,0.9,1.1$, respectively. The results are presented in Table 2.
@@
The behavior of the lensing coefficients $\bar{a}$ and $\bar{b}$ are expressed as functions of the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$ (see Figs. 5 and 6). It is observed from Fig. 3 that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase while keeping the other parameters fixed, the lensing coefficient $\bar{a}$ increases. Similarly, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters held constant, also results in a decrease in the value of the lensing coefficient $\bar{a}$ (see Table 2). The lensing coefficient $\bar{a}$ reaches the minimum value $\bar{a}=1$ [56] when both $S_{0}\rightarrow 0$ and $S_{2}\rightarrow 0$, corresponding to Schwarzschild black hole. Furthermore, Fig. 6 shows that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase, while keeping the other parameters fixed, the lensing coefficient $\bar{b}$ first decreases, reaches a minimum value, and then increases. Conversely, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters held constant, results in an increase in the value of the lensing coefficient $\bar{b}$ (see Table 2). The lensing coefficient $\bar{b}$ reaches the value $\bar{b}=-0.40023$ [56] when both $S_{0}\rightarrow 0$ and $S_{2}\rightarrow 0$, corresponding to Schwarzschild black hole.
@@
The behavior of the strong deflection angle, $\alpha_{D}$, is expressed as a function of the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$ (see Fig. 7). It is observed from Fig. 7 that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase while keeping the other parameters fixed, the lensing coefficient $\bar{a}$ increases. Similarly, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters held constant, decreases the value of the strong deflection angle, $\alpha_{D}$. Fig. 7 also shows that $\alpha_{D}\rightarrow\infty$ as $u\rightarrow u_{ph}$, i.e., the strong deflection angle $\alpha_{D}$ unboundedly large when $u\rightarrow u_{ph}$. Furthermore, the strong deflection angle, $\alpha_{D}$, in the context of an asymptotically safe quantum gravity black hole is more significant than that of a Schwarzschild black hole. The asymptotically safe quantum gravity black hole, with the presence of parameters $S_{0}$, $S_{2}$, $m_{0}$, and $m_{2}$, can significantly intensify the gravitational lensing effect compared to other black holes. This result indicates that the gravitational lensing effect produced by an asymptotically safe quantum gravity black hole is significantly enhanced compared to ordinary astrophysical black holes such as the Schwarzschild black hole ($S_{0}=S_{2}=0$). Thus, the asymptotically safe quantum gravity black hole, influenced by moving fluids and sound waves, can be detected more quickly and distinguished from ordinary astrophysical black holes, such as the Schwarzschild black hole.
@@
Table 2: Estimation of the strong lensing coefficients for different values of the asymptotically safe quantum gravity black hole parameters $S_{0}=0,0.1$, $S_{2}=0,0.1,0.2,0.3$, $m_{0}=0.5,1$ and $m_{2}=0.3,0.6,0.9,1.1$ respectively.
@@ IV.2 Lensing observables
We now turn to exploring the observable features of strong gravitational lensing within the framework of asymptotically safe quantum gravity black hole geometry. In this analysis, we assume that both the observer and the source are positioned at large distances from the black hole (acting as the lens) with near-perfect alignment. Furthermore, the source is assumed to lie directly behind the black hole. Under these conditions, the lens equation is expressed as [121]
@@
where $\Delta\alpha_{n}=\alpha_{D}(\theta)-2n\pi$ denotes the offset deflection angle, with $\alpha_{D}(\theta)$ representing the primary deflection angle, and $n$ indicating the number of loops or revolutions a photon completes around the black hole. The angles $\beta$ and $\theta$ correspond to the angular separations between the lens (black hole) and the source and between the image and the source, respectively. The distances $D_{ol}$, $D_{ls}$, and $D_{os}$ refer to the observer-lens, lens-source, and observer-source distances, respectively, with the relation $D_{os}=D_{ol}+D_{ls}$.
@@
The above equation indicates that the first relativistic image exhibits the highest brightness, with the magnification decreasing exponentially as $n$ increases. In other words, the brightness of the first image exceeds that of the subsequent relativistic images. Moreover, Eq. (27) becomes unbounded as $\beta\rightarrow 0$, implying that perfect alignment enhances the chances of detecting relativistic images.
We focus on the scenario where the brightest image, i.e., the outermost image at $\theta_{1}$, is resolved as a distinct image, while the remaining inner images cluster near $\theta_{\infty}$, defined as $\theta_{n}|{n\rightarrow\infty}:=\theta{\infty}$. Utilizing the deflection angle from Eq. (22), one can derive key strong lensing observables, including the angular position of the image cluster at $\theta_{\infty}$, the angular separation $S$ between the outermost and innermost images, and the relative magnification $r_{mag}$ between the outermost and inner relativistic images. These quantities are given as [56, 64]:
@@
If the strong lensing observables $\theta_{\infty}$, $X$, and $r_{mag}$ are determined from observations, the lensing coefficients $\bar{a}$, $\bar{b}$, and the minimum impact parameter $u_{ph}$ can be computed by inverting Eqs. (29), (30), and (31), respectively. Additionally, the observed data can be compared with the corresponding theoretical predictions. These results enable the identification and distinction of the asymptotically safe quantum gravity black hole from a Schwarzschild black hole.
@@
Table 3: Estimation of the strong lensing observables for the Supermassive Black Holes $M87^{*}$, $SgrA^{*}$for different values of the asymptotically safe quantum gravity black hole parameters $S_{0}=0.1$, $m_{0}=0.5,1$, $m_{2}=0.3,0.6,0.9,1.1$ and $S_{2}=0.1,0.2,0.3$ respectively. The observable quantity $r_{mag}$ does not depend on the mass or the distance of the black hole from the observer.
IV.2.1 Lensing observables and the asymptotically safe quantum gravity black hole parameters
Considering the supermassive black holes $M87^{*}$ and $SgrA^{*}$ in the center of the nearby galaxies, we numerically obtain the observable quantities $\theta_{\infty}$ ,$X$ and $r_{mag}$ for the asymptotically safe quantum gravity black hole (see Table 3). The mass and distance from the earth for $M87^{*}$ [5] are $M\approx 6.5\times 10^{9}M_{\odot}$,$D_{ol}\approx 16.8Mpc$, for $SgrA^{*}$ are $M\approx 4.28\times 10^{6}M_{\odot}$,$D_{ol}\approx 8.32kpc$ [122] . The behavior of the observable quantities $\theta_{\infty}$ ,$X$ and $r_{mag}$ for the supermassive black holes $M87^{*}$ and $SgrA^{*}$ are expressed as functions of the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$ (see Figs. 8,9 and 10). It is observed from Fig. 8 that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase while keeping the other parameters fixed, the observable quantity $\theta_{\infty}$ decreases. Similarly, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters held constant, also increases the value of the observable quantity $\theta_{\infty}$ (see Table 3). The observable quantity $\theta_{\infty}$ reaches the maximum value when both $S_{0}\rightarrow 0$ and $S_{2}\rightarrow 0$, corresponding to Schwarzschild black hole. Furthermore, Fig. 9 shows that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase, while keeping the other parameters fixed, $X$ increases. Conversely, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters held constant, results in a decrease in the value of the observable quantity $X$ (see Table 3). The observable quantity $X$ reaches the minimum value when both $S_{0}\rightarrow 0$ and $S_{2}\rightarrow 0$, corresponding to Schwarzschild black hole.
@@ IV.3 Einstein rings
When the source, the black hole (lens), and the observer are perfectly aligned, corresponding to $\beta=0$, the black hole deflects light rays in all directions, producing a ring-shaped image. This phenomenon, known as the Einstein ring, has been extensively studied in [47, 123, 124, 125, 126, 127]. By simplifying Eq. (27) for $\beta=0$, the angular radius of the $n^{\text{th}}$ relativistic image can be expressed as
@@
In the scenario where the black hole (lens) is positioned at a distance halfway between the source and the observer, denoted as $D_{os}=2D_{ol}$, and assuming that $D_{ol}\gg u_{ph}$, the angular radius of the $n^{\text{th}}$ relativistic Einstein ring within the spacetime of an asymptotically safe quantum gravity black hole can be expressed as follows
@@
Figure 11: The outermost Einstein ring $\theta^{E}_{1}$ ($\mu as$) of the black hole in asymptotically safe quantum gravity, in the context of $M87^{*}$ and $SgrA^{*}$, is presented in the X-Y plane. The black hole parameters $S_{0}$ and $S_{2}$ are varied independently, while the other parameters remain fixed.
The angular radius $\theta^{E}_{1}$ represents the outermost Einstein ring, as displayed in Fig. 11 for the supermassive black holes $M87^{*}$ and $SgrA^{*}$. From Fig. 11, it is observed that the angular radius $\theta^{E}_{1}$ of the outermost Einstein ring decreases with the increasing value of the parameter $S_{0}$ or parameter $S_{2}$. From Fig. 11, it is also evident that the angular radius $\theta^{E}_{1}$ of the outermost Einstein ring for the asymptotically safe quantum gravity black hole is smaller than the corresponding angular radius of the Schwarzschild black hole.

### v3

@@ IV Strong gravitational lensing and its observables
In this section, we explore the phenomenon of strong gravitational lensing around a black hole in quadratic gravity. The aim is to investigate how the black hole parameters $S_{0}$, $S_{2}$, $m_{0}$, and $m_{2}$ affect the observable outcomes of strong lensing. Additionally, we compare the lensing behavior of the black hole in quadratic gravity with that of the Schwarzschild black hole ($S_{0}=0=S_{2}$).
We begin our analysis by examining the strong deflection angle of a light ray in the equatorial plane ($\theta=\frac{\pi}{2}$) of the black hole.
IV.1 The strong deflection angle
To obtain the strong deflection angle of a light ray in the equatorial plane ($\theta=\frac{\pi}{2}$) of a black hole, we rewrite first the metric (2) by redefining the quantities $r$, $t$, $S_{0}$, $S_{2}$, $m_{0}$ and $m_{2}$ in units of the radius $2M$, so that $t\rightarrow t/2M$, $r\rightarrow r/2M$, $S_{0}\rightarrow S_{0}/2M$,$S_{2}\rightarrow S_{2}/2M$,$m_{0}\rightarrow 2Mm_{0}$ and $m_{2}\rightarrow 2Mm_{2}$, respectively. Hence, the quadratic geometric metric (2) takes in the equatorial approximation the form
@@ IV.1 The strong deflection angle
The unstable circular photon orbit, characterized by the radius $r_{\text{ph}}$, can be determined by applying the conditions of the effective potential, given by $\left.\frac{dV_{eff}}{dr}\right|_{r_{ph}}=0$ and $\left.\frac{d^{2}V_{eff}}{dr^{2}}\right|_{r_{ph}}<0$. Thus, the radius of the photon sphere $r_{ph}$ is the largest real root of the equation [116, 117, 118]:
@@
The behavior of the photon sphere radius, $r_{ph}$ and the critical impact parameter, $u_{ph}$, are expressed as functions of the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$ (see Figs. 5 and 6). It is observed from Fig. 5 that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase while the other parameters are kept fixed, the photon sphere radius $r_{ph}$ decreases. In contrast, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters kept constant, increases the radius of the photon sphere, $r_{ph}$. Furthermore, the radius of the photon sphere $r_{ph}$ reaches its maximum ($r_{ph}=1.5$ ) when both $S_{0}\rightarrow 0$ and $S_{2}\rightarrow 0$, corresponding to the Schwarzschild black hole.
It is also observed in Fig. 6 that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase while the other parameters are kept fixed, the critical impact parameter, $u_{ph}$, decreases. In contrast, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters kept constant, results in an increase in the critical impact parameter, $u_{ph}$ (see Table 4). Furthermore, the critical impact parameter $u_{ph}$ reaches its maximum ($u_{ph}=2.59808$ ) when both $S_{0}\rightarrow 0$ and $S_{2}\rightarrow 0$, corresponding to Schwarzschild black hole.
@@
The strong deflection angle in the quadratic gravity black hole spacetime can be obtained as a function of the closest approach distance $r_{0}$ as [59]
@@
The strong deflection angle $\alpha_{D}(r_{0})$ depends on the relationship between $r_{0}$ and $r_{ph}$ and it increases when these two values are approximately equal (i.e., $r_{0}\approx r_{ph}$). Hence, we introduce a new variable $z$ defined as [119, 120]
@@
where the above integral is computed numerically. We numerically determine the values of the critical impact parameter $u_{\text{ph}}$ and the strong lensing coefficients $\bar{a}$ and $\bar{b}$ at $r=r_{\text{ph}}$ for the quadratic gravity black hole, using the following parameter values: $S_{0}=0,0.5$, $S_{2}=0,0.1,0.2,0.3$, $m_{0}=0.5,1$, and $m_{2}=0.3,0.6,0.9,1.1$. The results are presented in Table 4.
@@
The behavior of the lensing coefficients $\bar{a}$ and $\bar{b}$ is expressed as functions of the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$ (see Figs. 7 and 8). It is observed from Fig. 7 that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase while keeping the other parameters fixed, the lensing coefficient $\bar{a}$ increases. Similarly, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters kept constant, also results in a decrease in the value of the lensing coefficient $\bar{a}$ (see Table 4). The lensing coefficient $\bar{a}$ reaches the minimum value $\bar{a}=1$ [65] when both $S_{0}\rightarrow 0$ and $S_{2}\rightarrow 0$, corresponding to the Schwarzschild black hole. Furthermore, Fig. 8 shows that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase, while the other parameters are kept fixed, the lensing coefficient $\bar{b}$ first decreases, reaches a minimum value and then increases. In contrast, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters kept constant, results in an increase in the value of the lensing coefficient $\bar{b}$ (see Table 4). The lensing coefficient $\bar{b}$ reaches the value $\bar{b}=-0.40023$ [65] when both $S_{0}\rightarrow 0$ and $S_{2}\rightarrow 0$, corresponding to the Schwarzschild black hole.
@@
The behavior of the strong deflection angle, $\alpha_{D}$, is expressed as a function of the pair of parameters $(S_{0},S_{2})$ and $(m_{0},m_{2})$ (see Fig. 9). It is observed from Fig. 9 that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase while keeping the other parameters fixed, the strong deflection angle $\alpha_{D}$ increases. Similarly, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters held constant, decreases the value of the strong deflection angle, $\alpha_{D}$. Fig. 9 also shows that $\alpha_{D}\rightarrow\infty$ as $u\rightarrow u_{ph}$, that is, the strong deflection angle $\alpha_{D}$ is unboundedly large when $u\rightarrow u_{ph}$. Furthermore, the strong deflection angle, $\alpha_{D}$, in the context of a quadratic gravity black hole is more significant than that of a Schwarzschild black hole. The quadratic gravity black hole, with the presence of parameters $S_{0}$, $S_{2}$, $m_{0}$ and $m_{2}$, can significantly intensify the gravitational lensing effect compared to other black holes. This result indicates that the gravitational lensing effect produced by a quadratic gravity black hole is significantly enhanced compared to ordinary astrophysical black holes such as the Schwarzschild black hole ($S_{0}=S_{2}=0$). Thus, the quadratic gravity black hole, influenced by moving fluids and sound waves, can be detected more quickly and distinguished from ordinary astrophysical black holes, such as the Schwarzschild black hole.
@@
Table 4: Estimation of the strong lensing coefficients for different values of the quadratic gravity black hole parameters $S_{0}=0,0.1$, $S_{2}=0,0.1,0.2,0.3$, $m_{0}=0.5,1$ and $m_{2}=0.3,0.6,0.9,1.1$ respectively.
@@ IV.2 Lensing observables
We now turn to exploring the observable features of strong gravitational lensing within the framework of quadratic gravity black hole geometry. In this analysis, we assume that both the observer and the source are positioned at large distances from the black hole (acting as the lens) with near-perfect alignment. Furthermore, the source is assumed to lie directly behind the black hole. Under these conditions, the lens equation is expressed as [122]
@@
where $\Delta\alpha_{n}=\alpha_{D}(\theta)-2n\pi$ denotes the offset deflection angle, with $\alpha_{D}(\theta)$ representing the primary deflection angle, and $n$ indicating the number of loops or revolutions a photon completes around the black hole. The angles $\beta$ and $\theta$ correspond to the angular separations between the lens (black hole) and the source and between the image and the source, respectively. The distances $D_{ol}$, $D_{ls}$ and $D_{os}$ refer to the observer-lens, lens-source and observer-source distances, respectively, with the relation $D_{os}=D_{ol}+D_{ls}$.
@@
The above equation indicates that the first relativistic image exhibits the highest brightness, with the magnification decreasing exponentially as $n$ increases. In other words, the brightness of the first image exceeds that of the subsequent relativistic images. Moreover, Eq. (29) becomes unbounded as $\beta\rightarrow 0$, implying that perfect alignment improves the chances of detecting relativistic images.
We focus on the scenario where the brightest image, i.e., the outermost image at $\theta_{1}$, is resolved as a distinct image, while the remaining inner images cluster near $\theta_{\infty}$, defined as $\theta_{n}|{n\rightarrow\infty}:=\theta{\infty}$. Using the deflection angle from Eq. (23), one can derive key strong lensing observables, including the angular position of the image cluster at $\theta_{\infty}$, the angular separation $S$ between the outermost and innermost images, and the relative magnification $r_{mag}$ between the outermost and inner relativistic images. These quantities are given as [65, 73]:
@@
If the strong lensing observables $\theta_{\infty}$, $X$ and $r_{mag}$ are determined from observations, the lensing coefficients $\bar{a}$, $\bar{b}$ and the minimum impact parameter $u_{ph}$ can be computed by inverting Eqs. (30), (31) and (32), respectively. Additionally, the observed data can be compared with the corresponding theoretical predictions. These results enable the identification and distinction of the quadratic gravity black hole from a Schwarzschild black hole.
@@
Table 5: Estimation of the strong lensing observables for the Supermassive Black Holes $M87^{*}$, $SgrA^{*}$ for different values of the quadratic gravity black hole parameters $S_{0}=0.1$, $m_{0}=0.5,1$, $m_{2}=0.3,0.6,0.9,1.1$ and $S_{2}=0.1,0.2,0.3$ respectively. The observable quantity $r_{mag}$ does not depend on the mass or the distance of the black hole from the observer.
IV.2.1 Lensing observables and the quadratic gravity black hole parameters
Considering the supermassive black holes $M87^{*}$ and $SgrA^{*}$ in the center of the nearby galaxies, we numerically obtain the observable quantities $\theta_{\infty}$ ,$X$ and $r_{mag}$ for the quadratic gravity black hole (see Table 5). The mass and distance from the earth for $M87^{*}$ [5] are $M\approx 6.5\times 10^{9}M_{\odot}$,$D_{ol}\approx 16.8Mpc$, for $SgrA^{*}$ are $M\approx 4.28\times 10^{6}M_{\odot}$,$D_{ol}\approx 8.32kpc$ [123] . The behavior of the observable quantities $\theta_{\infty}$ ,$X$ and $r_{mag}$ for the supermassive black holes $M87^{*}$ and $SgrA^{*}$ are expressed as functions of the parameter pairs $(S_{0},S_{2})$ and $(m_{0},m_{2})$ (see Figs. 10,11 and 12). It is observed from Fig. 10 that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase while keeping the other parameters fixed, the observable quantity $\theta_{\infty}$ decreases. Similarly, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters held constant, also increases the value of the observable quantity $\theta_{\infty}$ (see Table 5). The observable quantity $\theta_{\infty}$ reaches the maximum value when both $S_{0}\rightarrow 0$ and $S_{2}\rightarrow 0$, corresponding to Schwarzschild black hole. Furthermore, Fig. 11 shows that as the numerical values of the parameters $S_{0}>0$ or $S_{2}>0$ increase, while keeping the other parameters fixed, $X$ increases. Conversely, an increase in the values of the parameters $m_{0}>0$ or $m_{2}>0$, with the other parameters held constant, results in a decrease in the value of the observable quantity $X$ (see Table 5). The observable quantity $X$ reaches the minimum value when both $S_{0}\rightarrow 0$ and $S_{2}\rightarrow 0$, corresponding to Schwarzschild black hole.
@@ IV.3 Einstein rings
When the black hole (lens), the source, and the observer are perfectly aligned, corresponding to $\beta=0$, the black hole bends light in every direction, creating a ring-like image. This phenomenon, known as the Einstein ring, has been extensively studied in [56, 124, 125, 126, 127, 128]. By simplifying Eq. (28) for $\beta=0$, the angular radius of the $n^{\text{th}}$ relativistic image can be expressed as
@@
In the case where the black hole (lens) is situated halfway between the source and the observer, with $D_{os}=2D_{ol}$, and assuming that $D_{ol}\gg u_{ph}$, the angular radius of the $n^{\text{th}}$ relativistic Einstein ring within the spacetime of a quadratic gravity black hole can be expressed as follows
@@
Figure 13: The outermost Einstein ring $\theta^{E}_{1}$ ($\mu as$) of the black hole in quadratic gravity, in the context of $M87^{*}$ and $SgrA^{*}$, is presented in the X-Y plane. The black hole parameters $S_{0}$ and $S_{2}$ are varied independently, while the other parameters remain fixed.
The angular radius $\theta^{E}_{1}$ represents the outermost Einstein ring, as shown in Fig. 13 for supermassive black holes $M87^{*}$ and $SgrA^{*}$. From Fig. 13, it is observed that the angular radius $\theta^{E}_{1}$ of the outermost Einstein ring decreases with increasing value of the parameter $S_{0}$ or parameter $S_{2}$. From Fig. 13, it is also evident that the angular radius $\theta^{E}_{1}$ of the outermost Einstein ring for the quadratic gravity black hole is smaller than the corresponding angular radius of the Schwarzschild black hole.